OpenStudy (anonymous):

Perform the indicated operations and simplify. (-6x^2y-10xy^2)(-8x^2y^3-4y^2)

OpenStudy (gw2011):

Multiplying each member you get: (-6x^2y)(-8x^2y^3)=48x^4y^4 (-6x^2y)(-4y^2)=24x^2y^3 (-10xy^2)(-8x^2y^3)=80x^3y^5 (-10xy^2)(-4y^2)=40xy^4 All this comes out to: 48x^4y^4+24x^2y^3+80x^3y^5+40xy^4 Factoring this out you get: 8xy^3(6x^3y+3x+10x^2y^2+5y)